हिंदी

Six new employees, two of whom are married to each other, are to be assigned six desks that are lined up in a row. If the assignment of employees to desks is made randoml

Advertisements
Advertisements

प्रश्न

Six new employees, two of whom are married to each other, are to be assigned six desks that are lined up in a row. If the assignment of employees to desks is made randomly, what is the probability that the married couple will have nonadjacent desks?

योग
Advertisements

उत्तर

Number of desk occupied by one couple = 1

Only (4 + 1) = 5 persons to be assigned.

∴ Number of ways of assigning these 5 persons = 5! × 2!

Total number of ways of assigning 6 persons = 6!

∴ Probability that a couple has adjacent desk = `(5! xx 2!)/(6!) = 1/3`

So, the probability that the married couple will have no-adjacent desks = `1 - 1/3 = 2/3`.

Hence, the required probability = `2/3`.

shaalaa.com
  क्या इस प्रश्न या उत्तर में कोई त्रुटि है?
अध्याय 16: Probability - Exercise [पृष्ठ २९६]

APPEARS IN

एनसीईआरटी एक्झांप्लर Mathematics Exemplar [English] Class 11
अध्याय 16 Probability
Exercise | Q 2 | पृष्ठ २९६

वीडियो ट्यूटोरियलVIEW ALL [1]

संबंधित प्रश्न

Which of the following can not be valid assignment of probabilities for outcomes of sample space S = {ω1, ω2,ω3,ω4,ω5,ω6,ω7}

Assignment ω1 ω2 ω3 ω4 ω5 ω6 ω7
(a) 0.1 0.01 0.05 0.03 0.01 0.2 0.6
(b) `1/7` `1/7` `1/7` `1/7` `1/7` `1/7` `1/7`
(c) 0.1 0.2 0.3 0.4 0.5 0.6 0.7
(d) –0.1 0.2 0.3 0.4 -0.2 0.1 0.3
(e) `1/14` `2/14` `3/14` `4/14` `5/14` `6/14` `15/14`

A coin is tossed twice, what is the probability that at least one tail occurs?


A fair coin with 1 marked on one face and 6 on the other and a fair die are both tossed. Find the probability that the sum of numbers that turn up is (i) 3 (ii) 12


In a lottery, person chooses six different natural numbers at random from 1 to 20, and if these six numbers match with the six numbers already fixed by the lottery committee, he wins the prize. What is the probability of winning the prize in the game? [Hint: order of the numbers is not important.]


Check whether the following probabilities P(A) and P(B) are consistently defined

P(A) = 0.5, P(B) = 0.7, P(A ∩ B) = 0.6


Fill in the blank in following table:

P(A) P(B) P(A ∩ B) P(A ∪ B)
0.35 ... 0.25 0.6

4 cards are drawn from a well-shuffled deck of 52 cards. What is the probability of obtaining 3 diamonds and one spade?


Two unbiased dice are thrown. Find the probability that the sum of the numbers obtained on the two dice is neither a multiple of 2 nor a multiple of 3


A bag contains 8 red, 3 white and 9 blue balls. If three balls are drawn at random, determine the probability that all the three balls are blue balls 


A bag contains 8 red, 3 white and 9 blue balls. If three balls are drawn at random, determine the probability that  all the balls are of different colours.


If a letter is chosen at random from the English alphabet, find the probability that the letter is a consonant .


Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S =  {w1, w2, w3, w4, w5, w6, w7}:

Elementary events: w1 w2 w3 w4 w5 w6 w7
(iv)
\[\frac{1}{14}\]
\[\frac{2}{14}\]
\[\frac{3}{14}\]
\[\frac{4}{14}\]
\[\frac{5}{14}\]
\[\frac{6}{14}\]
\[\frac{15}{14}\]

A box contains 100 bulbs, 20 of which are defective. 10 bulbs are selected for inspection. Find the probability that: all 10 are defective


A box contains 100 bulbs, 20 of which are defective. 10 bulbs are selected for inspection. Find the probability that all 10 are good


Two dice are thrown together. The probability that neither they show equal digits nor the sum of their digits is 9 will be


An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is blue or white


An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is numbered 1, 2, 3, 4 or 5


An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is red or yellow and numbered 1, 2, 3 or 4


An urn contains twenty white slips of paper numbered from 1 through 20, ten red slips of paper numbered from 1 through 10, forty yellow slips of paper numbered from 1 through 40, and ten blue slips of paper numbered from 1 through 10. If these 80 slips of paper are thoroughly shuffled so that each slip has the same probability of being drawn. Find the probabilities of drawing a slip of paper that is numbered 5, 15, 25, or 35


One mapping (function) is selected at random from all the mappings of the set A = {1, 2, 3, ..., n} into itself. The probability that the mapping selected is one to one is ______.


If the letters of the word ALGORITHM are arranged at random in a row what is the probability the letters GOR must remain together as a unit?


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine P(A)


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine `P(A ∩ barB)`


A bag contains 8 red and 5 white balls. Three balls are drawn at random. Find the probability that all the three balls are white


If the letters of the word ASSASSINATION are arranged at random. Find the probability that four S’s come consecutively in the word


The probability that a student will pass his examination is 0.73, the probability of the student getting a compartment is 0.13, and the probability that the student will either pass or get compartment is 0.96.


If A and B are two candidates seeking admission in an engineering College. The probability that A is selected is .5 and the probability that both A and B are selected is at most .3. Is it possible that the probability of B getting selected is 0.7?


C1
Probability
C2
Written Description
(a) 0.95 (i) An incorrect assignment
(b) 0.02 (ii) No chance of happening
(c) – 0.3 (iii) As much chance of happening as not
(d) 0.5 (iv) Very likely to happen
(e) 0 (v) Very little chance of happening

A box contains 10 red marbles, 20 blue marbles and 30 green marbles. 5 marbles are drawn from the box, what is the probability that atleast one will be green?


If 4-digit numbers greater than 5,000 are randomly formed from the digits 0, 1, 3, 5, and 7, what is the probability of forming a number divisible by 5 when, the repetition of digits is not allowed?


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×